# Chip-Scale Point-Source Sagnac Interferometer by Phase-Space Squeezing

### v2

@@ Results
where $\Omega_{R}$ is the local Rabi frequency, $k_{0}=2\pi/\lambda_{0}$ is the optical transition wave-vector, $\Gamma$ is the spontaneous emission rate, $I(\mathbf{r})$ is the light intensity, and $c$ is the speed of light. The inverse harmonic repulsive potential could be implemented by the quadratic intensity profile near the center of a Gaussian beam propagating along the $z$ direction. However, a more efficient acceleration may be achieved by designing a fully quadratic beam shape. Here we consider acceleration in the $x$ direction induced by a laser beam propagating along the $y$ direction and having a homogeneous profile in the $z$ direction in the volume containing the atom cloud. The light beam profile is $I(x,z)=I_{0}(1-x^{2}/x_{0}^{2})$ for $|x|<x_{0}$, $|z|<z_{0}/2$ and zero otherwise, where $I_{0}$ is the peak intensity and $2x_{0}z_{0}$ is the beam cross-section. The peak intensity equals $I_{0}=3P/4x_{0}z_{0}$, where $P$ is the beam power, and the potential frequency becomes
$\omega_{\text{Harmonic}}=\sqrt{\frac{9\pi\Gamma P}{4mk_{0}^{3}c\Delta x_{0}^{3}z_{0}}}.$ (11)
@@
The optical repulsive potential has the consequent effect of heating the atoms due to scattered photons  (?). Along the direction of beam propagation, the atoms gain momentum in correlation to the spatial intensity $I(\textbf{r})$, which can be compensated by shifting the initial cloud position or, if necessary, by employing a counter-propagating beam. In addition, the scattering induces a random walk in velocity space in all directions, potentially enlarging the effective initial cloud size, $\tilde{\sigma}_{x0}$, and consequently decreasing the upper detection limit $\Omega_{\text{max}}$ (Eq. 6). Under the limitations we have taken for the laser power and typical parameter values, this effect reduces $\Omega_{\text{max}}$, e.g., by $\sqrt{2}$ for the parameters of Fig. 2. This effect can be reduced to a negligible level by improving the beam parameters, for example, by equally increasing the beam power $P$ and the detuning $\Delta$. More details about the heating process can be found in the appendix. One could also explore using different types of potential, such as magnetic gradients, to mitigate heating.
@@
Figure 3: Relative sensitivity and measured angular velocity deviation. (A) Relative sensitivity $\delta\Omega/\Omega$ vs. $\Omega$, where $\Omega$ is the nominal (actual) angular velocity, for the same simulation parameters and notation as in Fig. 2. The analytical curves are determined according to Eq. 5. The data points for PSI are not shown at large angular velocities since $\delta\Omega/\Omega$ exceeds the value of 1, thus irrelevant. (B) Simulation results of the measured angular velocity deviation $\Delta\Omega\equiv(\Omega_{\rm meas}-\Omega)$ for different angular velocities. Here $\Delta\Omega$ denotes the discrepancy between the measured and nominal angular velocities, while the sensitivity (uncertainty) $\delta\Omega$ is represented by the error bars. The data point labels are the same as in (A), while the black dashed line illustrates the point-source (PS) limit of zero deviation. The data points for PSI are not shown at large angular velocities since $\Delta\Omega$ is too large. The enhanced performance of the SPSI in dynamic range and sensitivity (measurement uncertainty) is clearly visible.
@@
In Fig. 2(A), the fringe pattern’s contrast is plotted against the angular velocity for both methods. This illustrates that SPSI can measure higher angular velocities as its contrast diminishes at higher angular velocities, much larger than for standard PSI. We briefly note that at large angular velocities, the fringe spatial frequency becomes high and the detection pixel resolution becomes a limiting factor, causing the numerical contrast to decay more rapidly than the analytically predicted decay, which does not account for this limitation. Similarly, the SPSI can measure lower angular velocities than the PSI because its final cloud radius is larger, rendering it more sensitive to slow angular velocities characterized by low spatial fringe frequency. This improvement in both limits is depicted in Fig. 2(B) by the analytical curve of the detection range as a function of $\eta$ according to Eq. 6.
@@
Fig. 3(A) presents a comparison of the relative sensitivity $\delta\Omega/\Omega$ between the two methods. It demonstrates an improvement of over one order of magnitude when both methods operate within their dynamic range $(60\lesssim\Omega\lesssim 200\,\rm{mrad/s})$, with a notably superior ratio beyond that range. Additionally, in Fig. 3(B), the plot of angular velocity deviation $\Delta\Omega=(\Omega_{\rm meas}-\Omega)$ provides further support for the aforementioned claims.
@@
To demonstrate the advantage of the SPSI in the compactness of the interferometer, we define the dimensionless parameter $a_{t}=T_{R,SPSI}/T_{R,PSI}$, representing the reduction in the Ramsey time in a compact SPSI relative to a standard PSI ($a_{t}<1$). Such a reduction would also reduce the fall distance $h$ during the interferometer by $a_{t}^{2}$. Additionally, the final cloud’s radius scales as $\sigma_{f}\sim a_{t}T_{\rm{ex}}\sigma_{v0}\eta$. Consequently, the sensitivity per shot is given by:
@@
$\delta\Omega>\frac{1}{a_{t}T_{R}^{2}\cdot 2K_{\text{eff}}C\sqrt{N/2}\cdot a_{t}\sigma_{v0}\eta},$ (12)
@@
such that the sensitivity per shot ratios become $\delta\Omega_{SPSI}/\delta\Omega_{PSI}=(C_{PSI}/C_{SPSI})/(a_{t}^{2}\cdot\eta$). For the design of a rotation sensor, it would be necessary to compromise between sensitivity and compactness and to identify the optimal operational point based on the parameters $\eta$ and $a_{t}$. Fig. 4 illustrates the advantage in single shot sensitivity offered by the SPSI over the standard PSI for different parameter values. It is evident from the figure that the incorporation of a repulsive potential allows for a sensitivity enhancement of up to two orders of magnitude, while concurrently reducing the cycle time by a factor of approximately two, or alternatively, decreasing the cycle time by around tenfold without compromising sensitivity.
@@
Figure 4: Sensitivity-compactness trade-off in the SPSI. The single shot sensitivity ratio $\delta\Omega_{SPSI}/\delta\Omega_{PSI}$ based on Eq. 12, is plotted against the squeezing parameter $\eta$ and the cycle time ratio $a_{t}=T_{SPSI}/T_{PSI}$ for an angular velocity of $\Omega=100\,\rm{mrad/s}$, while maintaining constant values for the remaining parameters: Time between pulses $T_{R}=10\,\rm{ms}$, temperature $T=2\,\rm{\mu K}$, and initial cloud size $\sigma_{x0}=100\,\rm{\mu m}$. The maximum value of $\eta$ was chosen to ensure that the majority of atoms are affected by the same potential throughout the acceleration phase. The bold black line indicates the validity boundary of this plot beyond which the minimal detectable angular velocity of the SPSI method (Eq. 13) exceeds the nominal angular velocity $\Omega$. It is evident that the implementation of a repulsive potential allows for a sensitivity enhancement of more than an order of magnitude, while simultaneously reducing the cycle time by a factor of about two (as indicated by the bright red color), or reducing the cycle time by approximately tenfold without affecting the sensitivity (as indicated by the light-blue color). Specifically, utilizing the compactness factor definition of Eq. 14, we find that with $\eta=40$ and $a_{t}=0.1$, we have a performance enhancement of $CF_{SPSI}/CF_{PSI}\simeq 2\cdot 10^{4}$ (black point).
@@
Specifically, we consider the following chip-scale scenario: a rectangular vacuum cell with dimensions of $1\,\rm{mm}\times 10\,\rm{mm}\times 10\,\rm{mm}$, with the short dimension aligned perpendicular to the direction of the repulsive potential’s acceleration. The SPSI sequence commences with the capture of an atomic cloud comprising $N=10^{6}$ 87Rb atoms, with a radius of $\sigma_{x0}=100\,\rm{\mu m}$ and temperature of $T=2\,\rm{\mu K}$, using a magneto-optical trap and optical molasses setup. Employing a fully harmonic potential characterized by dimensions of $x_{0}=200\,\rm{\mu m}$ and $z_{0}=400\,\rm{\mu m}$, with total power of $P=1\,\rm{W}$ and an acceleration time of $t_{\rm acc}=0.1\,\rm{ms}$, yields a squeezing parameter $\eta=46.2$, potential frequency $\omega_{\text{Harmonic}}=2\pi\cdot 1220\,$Hz, and final velocity uncertainty $\tilde{\sigma}_{v0}=0.64\,$m/s. Following acceleration, the atoms undergo a Ramsey sequence with a time $T_{R}=2.5\,\rm{ms}$, resulting in a total fall height of $h=120\,\rm{\mu m}$. The final cloud dimensions are $\sigma_{f,z}=170\,\rm{\mu m}$ height and $\sigma_{f,x}=0.33\,\rm{cm}$ width. Without heating, a vacuum dimension of $0.5\,\rm{mm}$ would suffice, but if heating is not mitigated, a vacuum dimension of $1\,\rm{mm}$ would be required, or alternatively, one would have to take into account that a considerable amount of atoms could not be recaptured for the next cycle, and reloading atoms from a source would be required.
With a number of operations per second of $\tau=100\,\rm{s^{-1}}$, the SPSI exhibits a sensitivity of $\delta\Omega=1\,\rm{\mu\rm{rad}/(s\cdot\sqrt{Hz})}$ and a one-shot sensitivity of $\delta\Omega=10\,\rm{\mu\rm{rad}/s}$. The minimum and maximum detectable angular velocities are $\Omega_{\rm min}=7.5\,\rm{mrad/s}$ and $\Omega_{\rm max}=11.47\,\rm{rad/s}$, respectively.
Studying the impact of heating due to the repulsive optical potential in a chip-scale device is crucial. The atoms acquire momentum in the direction of the optical repulsive potential beam, and this effect can be managed as discussed earlier. However, heating in the direction of the splitting and recombining laser beam presents more challenges, especially when we aim for the smallest possible chip. For the repulsive potential parameters outlined above, the calculation of the cloud’s final radius results in $\sigma_{z,f}\approx 300\,\mu\rm{m}$. This clearly remains within the defined $1\,\rm{mm}$ height limit. Additionally, as mentioned earlier, the dynamic range is restricted due to heating and is reduced in this example by a factor of approximately 3.
@@ Discussion
Analysis of the simulation results from Fig. 4 reveals that employing squeezing can lead to a tenfold reduction in the cycle time without compromising sensitivity. Conversely, sensitivity enhancement by up to two orders of magnitude is achievable while concurrently reducing the cycle time by about two-fold. Considering that sensitivity per unit time scales as the inverse square root of the repetition rate $\nu_{\text{p}}=1/\tau_{\text{rep}}$, and noting that the relation between the repetition rates of SPSI and standard PSI scales as $a_{t}$, we derive $\delta\Omega_{SPSI}/\delta\Omega_{PSI}=(C_{PSI}/C_{SPSI})/a_{t}^{1.5}\cdot\eta$.
@@
$CF\equiv\frac{\Omega_{\text{max}}/\Omega_{\text{min}}}{\delta\Omega\cdot h}.$ (14)
@@
This factor encapsulates sensitivity $\delta\Omega$, dynamic range (as the ratio $\Omega_{\rm max}/\Omega_{\rm min}$, and the vertical size of the interferometer $h$. Based on the previous comparison between SPSI and PSI, we observe an improvement ratio of $\eta^{3}\cdot\sqrt{a_{t}}$, which can exceed four orders of magnitude. For instance, considering the conditions depicted in Fig. 4, with $\eta=40$ and $a_{t}=0.1$, we derive $CF_{\rm SPSI}/CF_{\rm PSI}\simeq 2\cdot 10^{4}$.
@@
Finally, we note that details of the numerical analysis are presented in Materials and Methods, while the AC Stark interaction, heating by light and a full quantum approach are discussed in the appendices. As an outlook, let us note that accelerating the cloud expansion may also be realized by other interactions, such as repulsive magnetic forces or attractive electric forces. An accelerated expansion can even be acquired passively by diffracting a beam of atoms (e.g., from a 2D MOT) off a slit. An interesting problem left for future work is to find the optimal functional form of the repulsive potential.
@@
To conclude, the SPSI described in this work enables either enhanced performance for standard size devices or maintaining the performance while miniaturizing to a chip-scale device, opening the door to real-life applications. It may very well be that the squeezed source presented here may be adapted to serve other types of matter-wave interferometers and quantum technology devices.
@@
Figure 5: Simulation fringe pattern fit. This figure compares the simulation results of both methods, the top for PSI and the bottom for SPSI. The light blue shaded area depicts the total number of atoms in each bin, denoted as $N_{\text{total}}$, while the darker shade indicates only those atoms in the internal state $|1\rangle$, denoted as $N_{|1\rangle}$. The green points represent the normalized probability for the $|1\rangle$ state, calculated as $P_{1}=N_{|1\rangle}(x)/N_{\text{total}}(x)$, while the red line denotes the sinusoidal fit to these green points. The angular velocity is determined from the fit’s fringe spatial frequency using Eq. 3. The simulations begin with an initial 1D Gaussian cloud comprising $N=10^{6}$ atoms, with a width of $\sigma_{x0}=100\,\rm{\mu m}$ and a temperature of $T=5\,\rm{\mu K}$. For the PSI method, the atoms undergo free expansion during the interferometer sequence, with a time interval between pulses of $T_{R}=25\,\rm ms$, while the simulated angular velocity is $\Omega=40\,\rm{mrad/s}$. In the case of SPSI, the cloud is additionally subjected to a repulsive potential having a power of $P=1\,\rm{W}$, with a beam cross-section of $400\crossproduct 400\,\mu$m2, activated for $t_{\text{acc}}=20\,\rm{\mu s}$. This results in a potential frequency of $\omega=2\pi\times 1.214\,\rm{kHz}$ (Eq. 11), a squeezing parameter $\eta=5.4$ (Eq. 8), and an effective time $t_{\text{eff}}=0.84\,\rm{ms}$ (Eq. 9). Consequently, the cloud size increases notably. Note that the contrast and number of fringes in the PSI method is smaller than that of the SPSI (see Figs. 1-2), leading to a fit deviation and uncertainty one order of magnitude higher for the PSI case.

### v3

@@ Results
where $\Omega_{R}$ is the local Rabi frequency, $k_{0}=2\pi/\lambda_{0}$ is the optical transition wave-vector, $\Gamma$ is the spontaneous emission rate, $I(\mathbf{r})$ is the light intensity, and $c$ is the speed of light. The inverse harmonic repulsive potential could be implemented by the quadratic intensity profile near the centre of a Gaussian beam propagating along the $z$ direction. However, a more efficient acceleration may be achieved by designing a fully quadratic beam shape. For simplicity, we consider here a one-dimensional model of acceleration in the $x$ direction induced by a laser beam propagating along the $y$ direction and having a homogeneous profile in the $z$ direction in the volume containing the atom cloud. The light beam profile is $I(x,z)=I_{0}(1-4x^{2}/\Delta x_{\text{rep}}^{2})$ for $|x|<\Delta x_{\text{rep}}/2$, $|z|<\Delta z_{\text{rep}}/2$ and zero otherwise. Here the peak intensity $I_{0}$ is related to the beam power $P$ as $I_{0}=3P/2\Delta x_{\text{rep}}\Delta z_{\text{rep}}$. The harmonic frequency of the repulsive potential is then
$\omega_{\text{Harmonic}}=\sqrt{\frac{18\pi\Gamma P}{mk_{0}^{3}c\Delta\cdot\Delta x_{\text{rep}}^{3}\Delta z_{\text{rep}}}}.$ (13)
@@
The optical repulsive potential has the consequent effect of heating the atoms due to scattered photons  (?). Along the direction of beam propagation, the atoms gain momentum in correlation to the spatial intensity $I(\textbf{r})$, which can be compensated by shifting the initial cloud position or, if necessary, by employing a counter-propagating beam. In addition, the scattering induces a random walk in velocity space in all directions, potentially enlarging the effective initial cloud size, $\tilde{\sigma}_{x0}$, and consequently decreasing the upper detection limit $\Omega_{\text{max}}$ (Eq. 8). Under the limitations we have taken for the laser power and typical parameter values, this effect reduces $\Omega_{\text{max}}$, e.g., by $\sqrt{2}$ for the parameters of Fig. 2. This effect can be reduced to a negligible level by improving the beam parameters, for example, by equally increasing the beam power $P$ and the detuning $\Delta$. More details about the heating process can be found in the supplementary information. One could also explore using different types of potential, such as magnetic gradients, to mitigate heating.
Let us note that effective phase-space squeezing is also the natural result of the expansion of a Bose-Einstein condensate (BEC) with repulsive atom-atom interactions. A BEC that is initially trapped in a cigar-shaped potential expands along its fast axes, after turning off the trap, such that its size scales like $\sigma(t)\approx\sigma(0)\sqrt{1+\omega^{2}t^{2}}$, where $\omega$ is the trap frequency  (?). After a time $t\gg\omega^{-1}$ the momentum width of the expanding wave packet $\sigma_{p}\approx m\sigma(0)\omega$ corresponds to an equivalent wave packet with no atom-atom interactionswith an initial position uncertainty $\sigma_{x0}=\hbar/2\sigma_{p}$. This initial equivalent wave-packet size is smaller than the minimum uncertainty wave packet with $\sigma_{\rm min}=\sqrt{\hbar/2m\omega}$ by a factor $\eta_{BEC}=\sigma_{\rm min}/\sigma(0)<1$ and smaller than the actual initial BEC wave-packet by a factor $\eta_{BEC}^{2}$. This property may be used to create a collimated beam by the method of delta-kick cooling, but in our case it may be an advantage without this kick due to the fast expansion and high velocity-position correlation. However, in this paper we concentrate on an initial thermal atomic cloud without atom-atom interaction due to its practical advantages over a BEC in this context.
@@
Figure 3: Relative single-shot sensitivity and measured angular velocity deviation. (A) Relative sensitivity $\delta\Omega_{\text{single}}/\Omega$ vs $\Omega$, where $\Omega$ is the nominal angular velocity, for the same simulation parameters and notation as in Fig. 2. The analytical curves are determined according to Eq. 7. The simulation results for PSI at large angular velocities where $\delta\Omega_{\text{single}}/\Omega>1$ are irrelevant and therefore not shown. (B) Simulation results of the measured angular velocity deviation $\Delta\Omega$ (Eq. 6) for different angular velocities. Here $\Delta\Omega$ the discrepancy between the measured and nominal angular velocities, while the sensitivity (uncertainty) $\delta\Omega_{\text{single}}$ is represented by the error bars. $\Omega$ is extracted from each simulation run and obtained from the fringe spatial frequency of the fitted sinusoidal pattern (as described in Fig. 5). The data point labels are the same as in (A), while the black dashed line illustrates the point-source (PS) limit of zero deviation. The simulation results for PSI are not shown for large angular velocities where $\delta\Omega_{\rm single}/\Omega>1$. The enhanced performance of the SPSI in dynamic range and sensitivity (measurement uncertainty) is clearly visible.
@@
In Fig. 2(A), the fringe pattern’s contrast is plotted against the angular velocity for both methods. This illustrates that SPSI can measure higher angular velocities as its contrast diminishes at higher angular velocities, much larger than for standard PSI. We briefly note that at large angular velocities, the fringe spatial frequency becomes high, and the detection pixel resolution becomes a limiting factor, causing the numerical contrast to decay more rapidly than the analytically predicted decay, which does not account for this limitation. Similarly, the SPSI can measure lower angular velocities than the PSI because its final cloud radius is larger, rendering it more sensitive to slow angular velocities characterized by low spatial fringe frequency. This improvement in both limits is depicted in Fig. 2(B) by the analytical curve of the detection range as a function of $\eta$ according to Eq. 8.
@@
Fig. 3(A) presents a comparison of the relative sensitivity $\delta\Omega_{\text{single}}/\Omega$ between the two methods. It demonstrates an improvement of over one order of magnitude when both methods operate within their dynamic range $(60\lesssim\Omega\lesssim 200\,\rm{mrad/s})$, with a notably superior ratio beyond that range. Additionally, in Fig. 3(B), the plot of angular velocity deviation $\Delta\Omega$ (Eq. 6) provides further support for the advantage of SPSI.
@@
Having established the clear advantage of the SPSI when compared to PSI with equivalent interferometer durations, where the transverse size of the final cloud is much larger for the SPSI, we examine the SPSI’s advantages in terms of spatial compactness, particularly in scenarios with reduced operational timeframes where a conventional PSI would be nonfunctional due to an inadequate final cloud size relative to the initial cloud dimensions. To quantify the SPSI’s interferometric compactness, we introduce the dimensionless parameter $a_{t}=T_{R,SPSI}/T_{R,PSI}$, which represents the Ramsey time reduction in a compact SPSI compared to a standard PSI ($a_{t}<1$). This reduction would also reduce the fall distance $h$ during the interferometer by $a_{t}^{2}$ and the repetition rate $\nu$ by $a_{t}$. Additionally, the radius of the final cloud scales as $\sigma_{f}\sim a_{t}T_{\rm{ex}}\sigma_{v0}\eta$. Consequently, the sensitivity given by:
@@
$\delta\Omega\approx\frac{1}{a_{t}T_{R}^{2}\cdot 2K_{\text{eff}}C\sqrt{N\nu/2a_{t}}\cdot a_{t}\sigma_{v0}\eta},$ (14)
@@
such that the sensitivity per shot ratio become $\delta\Omega_{SPSI}/\delta\Omega_{PSI}=(C_{PSI}/C_{SPSI})/(a_{t}^{1.5}\cdot\eta$), where it was assumed that the interferometer starts right after the beginning of the expansion, such that $T_{\rm ex}\approx 2T_{R}$. For the design of a rotation sensor, it would be necessary to compromise between sensitivity and compactness and to identify the optimal operational point based on the parameters $\eta$ and $a_{t}$. Fig. 4 illustrates the advantage in sensitivity offered by the SPSI over the standard PSI for different parameter values. It is evident from the figure that the incorporation of a repulsive potential allows for a sensitivity enhancement of up to two orders of magnitude while concurrently reducing the cycle time by a factor of approximately two, or alternatively, decreasing the cycle time by around tenfold without compromising sensitivity.
@@
Figure 4: Sensitivity-compactness trade-off in the SPSI. The sensitivity ratio $\delta\Omega_{SPSI}/\delta\Omega_{PSI}$ based on Eq. 14, is colour coded against the squeezing parameter $\eta$ and the cycle time ratio $a_{t}=T_{SPSI}/T_{PSI}$ for an angular velocity of $\Omega=100\,\rm{mrad/s}$, while maintaining constant values for the remaining parameters: Time between pulses $T_{R}=10\,\rm{ms}$, temperature $T=2\,\rm{\mu K}$, and initial cloud size $\sigma_{x0}=100\,\rm{\mu m}$. The maximum value of $\eta$ was chosen to ensure that the majority of atoms are affected by the same potential throughout the acceleration phase. The bold black line indicates the validity boundary of this plot beyond which the minimal detectable angular velocity of the SPSI method (Eq. 15) exceeds the nominal angular velocity $\Omega$. It is evident that the implementation of a repulsive potential allows for a sensitivity enhancement of more than an order of magnitude while simultaneously reducing the cycle time by a factor of about two (as indicated by the bright red colour) or reducing the cycle time by approximately tenfold without affecting the sensitivity (as indicated by the light-blue colour). Specifically, utilizing the compactness factor definition of Eq. 16, we find that with $\eta=40$ and $a_{t}=0.1$, we have a performance enhancement of $\kappa_{SPSI}/\kappa_{PSI}\simeq 2\cdot 10^{4}$ (black point).
@@
Specifically, we consider the following chip-scale scenario: a rectangular vacuum cell with dimensions of $1\,\rm{mm}\times 10\,\rm{mm}\times 10\,\rm{mm}$, with the short dimension aligned perpendicular to the direction of the repulsive potential’s acceleration. The SPSI sequence commences with the capture of a thermal atomic cloud comprising $N=10^{6}$ 87Rb atoms, with a radius of $\sigma_{x0}=100\,\rm{\mu m}$ and temperature of $T=2\,\rm{\mu K}$, using a magneto-optical trap and optical molasses setup. Employing a fully harmonic potential characterized by dimensions of $\Delta x_{\text{rep}}=400\,\rm{\mu m}$ and $\Delta z_{\text{rep}}=400\,\rm{\mu m}$, with total power of $P=1\,\rm{W}$ and an acceleration time of $t_{\rm acc}=0.1\,\rm{ms}$, yields a squeezing parameter $\eta=46.2$, potential frequency $\omega_{\text{Harmonic}}=2\pi\cdot 1220\,$Hz, and final velocity uncertainty $\tilde{\sigma}_{v0}=0.64\,$m/s. Following acceleration, the atoms undergo a Ramsey sequence with a time $T_{R}=2.5\,\rm{ms}$, resulting in a total fall height of $h=120\,\rm{\mu m}$. The final cloud dimensions are $\sigma_{f,z}=170\,\rm{\mu m}$ height and $\sigma_{f,x}=0.33\,\rm{cm}$ width. Without heating, a vacuum dimension of $0.5\,\rm{mm}$ would suffice, but if heating is not mitigated, a vacuum dimension of $1\,\rm{mm}$ would be required, or alternatively, one would have to take into account that a considerable amount of atoms could not be recaptured for the next cycle, and reloading atoms from a source would be required.
With a number of operations per second $\nu=100\,\rm{s^{-1}}$, the SPSI exhibits a sensitivity of $\delta\Omega=1\,\rm{\mu\rm{rad}/(s\cdot\sqrt{Hz})}$ and a one-shot sensitivity of $\delta\Omega=10\,\rm{\mu\rm{rad}/s}$. The minimum and maximum detectable angular velocities are $\Omega_{\rm min}=7.5\,\rm{mrad/s}$ and $\Omega_{\rm max}=11.47\,\rm{rad/s}$, respectively.
Studying the impact of heating due to the repulsive optical potential in a chip-scale device is crucial. The atoms acquire momentum in the direction of the optical repulsive potential beam, and this effect can be managed, as discussed earlier. However, heating in the direction of the splitting and recombining laser beam presents more challenges, especially when we aim for the smallest possible chip. For the repulsive potential parameters outlined above, the calculation of the size of the cloud along the direction of splitting results in $\sigma_{z,f}\approx 300\,\mu\rm{m}$, such that most of the atoms remain within the $1\,\rm{mm}$ of the device. Additionally, as mentioned earlier, the dynamic range is restricted due to heating and is reduced in this example by a factor of approximately 3.
In a fully integrated chip-scale device it will be necessary to design photonics that will provide the shapes of the laser beams responsible for the repulsive potential and the stimulated Raman transitions such that the former has the quadratic intensity profile along $x$ and the latter has a flat intensity profile over the atomic cloud in both $x$ and $y$ directions. A beam that has a flat intensity across the significant spatial extent of the atomic cloud in $x$, can be generated, for example, using a set of beam splitters, as commonly implemented in 2D-MOT configurations  (?).
@@ Discussion
The presented method offers multiple orders of magnitude of improvement in the sensitivity and dynamic range of point-source interferometers. Our theoretical and numerical analysis demonstrates the potential of SPSI to significantly enhance performance in both standard-sized and chip-scale devices. The results highlight the substantial benefits of position-momentum phase-space squeezing, which include shorter cycle times and higher repetition rates. While our findings suggest that SPSI can pave the way for new practical applications, it is essential to further investigate the more minute and technical details of the method. Future experimental implementations should focus on optimizing parameters such as the repulsive potential, cloud size, Raman beam’s profile and detection techniques to fully realize the theoretical advantages of SPSI.
Analysis of the simulation results from Fig. 4 reveals that employing squeezing can lead to a tenfold reduction in the cycle time without compromising sensitivity. Conversely, sensitivity enhancement by up to two orders of magnitude is achievable while concurrently reducing the cycle time by about two-fold.
@@
$\kappa\equiv\frac{\Omega_{\text{max}}/\Omega_{\text{min}}}{\delta\Omega\cdot h}.$ (16)
@@
This factor encapsulates sensitivity $\delta\Omega$, dynamic range (as the ratio $\Omega_{\rm max}/\Omega_{\rm min}$, and the vertical size of the interferometer $h$. Based on the previous comparison between SPSI and PSI, we observe an improvement ratio of $\eta^{3}\cdot\sqrt{a_{t}}$, which can exceed four orders of magnitude. For instance, considering the conditions depicted in Fig. 4, with $\eta=40$ and $a_{t}=0.1$, we derive $\kappa_{\rm SPSI}/\kappa_{\rm PSI}\simeq 2\cdot 10^{4}$.
@@
Finally, we note that details of the numerical analysis are presented in Materials and Methods, while the AC Stark interaction, heating by light and a full quantum approach are discussed in the supplementary information. As an outlook, let us note that accelerating the cloud expansion may also be realized by other interactions, such as repulsive magnetic forces or attractive electric forces. An accelerated expansion can even be acquired passively by diffracting a beam of atoms (e.g., from a 2D MOT) off a slit. An interesting problem left for future work is to find the optimal functional form of the repulsive potential.
@@
To conclude, the SPSI described in this work enables either enhanced performance for standard-size devices or maintaining the performance while miniaturizing it to a chip-scale device, opening the door to real-life applications.
@@
Figure 5: Simulation fringe pattern fit. This figure compares the simulation results of both methods, the top for PSI and the bottom for SPSI. The light blue shaded area depicts the total number of atoms in each bin, denoted as $N_{\text{total}}$, while the darker shade indicates only those atoms in the internal state $|1\rangle$, denoted as $N_{|1\rangle}$. The green points represent the normalized probability for the $|1\rangle$ state, calculated as $P_{1}=N_{|1\rangle}(x)/N_{\text{total}}(x)$, while the red line denotes the sinusoidal fit to these green points. The angular velocity is determined from the fit’s fringe spatial frequency using Eq. 4. In this example the initial 1D Gaussian cloud comprising $N=10^{6}$ atoms, with a width of $\sigma_{x0}=100\,\rm{\mu m}$ and a temperature of $T=5\,\rm{\mu K}$, as in Figs. 2 and 3. For the PSI method’s simulation, the atoms undergo free expansion during the interferometer sequence, with a time interval between pulses of $T_{R}=25\,\rm ms$ used in this figure, while the simulated angular velocity is $\Omega=40\,\rm{mrad/s}$. In the case of SPSI, the cloud is additionally subjected to a repulsive potential having a power of $P=1\,\rm{W}$, with a beam cross-section of $400\crossproduct 400\,\mu$m2, as in Figs. 2 and 3, activated for $t_{\text{acc}}=20\,\rm{\mu s}$. This results in a potential frequency of $\omega=2\pi\times 1.214\,\rm{kHz}$ (Eq. 13), a squeezing parameter $\eta=5.4$ (Eq. 10), and an effective time $t_{\text{eff}}=0.84\,\rm{ms}$ (Eq. 11). Consequently, the cloud size increases notably, and we chose the potential parameters here such that it would be easy to compare to the PSI simulation. Note that the contrast and number of fringes in the PSI method is smaller than that of the SPSI (see Figs. 1-2), leading to a fit deviation and uncertainty one order of magnitude higher for the PSI case. It is important to note that while the histograms appear different in scale, the total atom count remains constant for both methods. The discrepancy in appearance is due to different bin sizes used in the PSI and SPSI plots to accommodate the varying cloud expansions, with SPSI utilizing larger bins to cover its greater spatial extent.